In a class of 40 students ,15 are boys.Find the percentage of students who are girls

Answer by stanbon(75887)

In a class of 40 students ,15 are boys.Find the percentage of students who are girls
---
# of girls = 40-15 = 25
----
Percent of girls = 25/40 = 5/8
Solve:: x/100 = 5/8
x = 500/8
x = 62.5
Ans: 62.5%
---------------------
Cheers,
Stan H.
----------------